OpenStudy (anonymous):

Please help:

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\sqrt{2x^2+9} dx\]

OpenStudy (anonymous):

find the value.

OpenStudy (lgbasallote):

have you tried using completing the square method?

OpenStudy (anonymous):

Looks like a candidate for trig substitution.

OpenStudy (anonymous):

i do know the basic radical, but integrating radicals like that?.....oh thats confusing....

OpenStudy (lgbasallote):

you should use completing the square first....that would give you an integral that looks like \[\int \sqrt{x^2 - a^2}dx\] does that look familiar?

OpenStudy (anonymous):

Thinks of the form \(\sqrt{a^2+b^2x^2} \) can be substituted with \(x=\frac{a}{b}\tan(\theta)\).

OpenStudy (anonymous):

\[a=3,\ b=\sqrt{2}\]See what happens...

OpenStudy (anonymous):

then what will be next?

OpenStudy (anonymous):

Well if you did the substitution, you'd get \[ x=\frac{3}{\sqrt{2}}\tan(\theta)\quad dx = \frac{3}{\sqrt{2}}\sec^2(\theta)d\theta \] \[ \Large \int_1^2 \sqrt{\sqrt{2}^2\left(\frac{3}{\sqrt{2}}\tan(\theta)\right)^2+3^2}\frac{3}{\sqrt{2}}\sec^2(\theta)d\theta \]

OpenStudy (anonymous):

Looks messy, but it really simplifies well.

OpenStudy (anonymous):

looks messy. then what will be next.....

OpenStudy (anonymous):

\[ \sqrt{\sqrt{2}^2\left(\frac{3}{\sqrt{2}}\tan(\theta)\right)^2+3^2} = \sqrt{3^2\tan^3(\theta)+3^2} = \sqrt{3^2(\tan^2(\theta)+1)} \]

OpenStudy (anonymous):

\[ = \sqrt{3^2\sec^2(\theta)} = 3\sec(\theta) \]

OpenStudy (anonymous):

what happened to \[\frac{ 3 }{ \sqrt{2} }\sec^2\theta d \theta\]?

OpenStudy (anonymous):

I was simplifying part of it. It's still there.

OpenStudy (anonymous):

ok......

OpenStudy (anonymous):

so how will i change it into \[\sec \theta d \theta\]